OpenStudy (anonymous):

solve 0.4x+4<=0.8x-9 my answer is x>=32.5 am i correct

OpenStudy (anonymous):

solve 5/4x+1/8x=9/8+x

OpenStudy (anonymous):

yes, you are right

OpenStudy (anonymous):

\[.4x + 4 \le .8x - 9 \]\[\implies -.4x \le -12\]\[\implies x \ge30\] Looks like you're a bit off.

OpenStudy (anonymous):

9+4 = 13, not 12 :)

OpenStudy (anonymous):

Doh! You're right. Foiled by arithmetic!

OpenStudy (anonymous):

lol :D